A and B are two thin concentric hollow conductors having radii a and 2 a and charge 2Q and Q respectively .if potential of outer sphere is 5V, then potential of inner sphere is

Grade:11

1 Answers

Shailendra Kumar Sharma
6 years ago
kQ/2a + 4kQ/2a)=4+4*5 =25VA and B are two thin concentric hollow conductors having radii a and 2 a and charge 2Q and Q respectively .if potential of outer sphere is 5V, then potential of inner sphereGiven potential at outer sphere kQ/2a =5;Potential in inner sphere is kQ/2a+k2Q/a=?